One number is 12 more than the other. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Found 2 solutions by checkley77, Mathtut:
Answer by checkley77(12844)

x & x+12 are the 2 numbers.
x+2(x+12)=39
x+2x+24=39
3x=39-24
3x=15
x=15/3
x=5 is the smaller number.
5+12=17 for the larger number.

Answer by Mathtut(3670)

lets call our two numbers a and b where a is the smaller and b is the larger. one number is 12 more than the other, so we write this as b=a+12 since b is the larger #. the sum of the smaller a+ twice the larger 2b equals 39 which we would write out as a+2b=39 now taking the a value in the first equation and substituting it into the second equation we get a+2(a+12)=39 which is the same as a+2a+24=39 now combine the a terms and subtract 24 on each side we get 3a=15 therefore a=5 and since b is 12 more than a or b=a+12 we get b=17
answer small number is 5
large number is 17